One number is five times another number. Their sum is twelve more than the smaller number. Find the two numbers.
Answer by stanbon(48568)

One number is five times another number. Their sum is twelve more than the smaller number. Find the two numbers.
---------------------------------------
Equations:
x = 5y
That means y is the smaller number.
x+y = y+12
-------------------
x = 12
y = (12/5)= 2.4
--------------------------
Cheers,
Stan H.